opdaxy:

The scale on a map is 2 cm = 25 miles. Dave measured the distance to the next town at 5 cm. How many miles away is the next town? @xiogonz

CourtneyWatson:

What do you think the answer is ?

opdaxy:

idk

XioGonz:

You need to occupy the simple rule of 3 in this case. Scale on the map--2cm=25m(miles) For each 2cm=real distance of 25m(miles) Here we see " Dave measured the distance to the next town at 5 cm." So 2cm-25m(miles) 5cm-xm(miles) 2x=25*5 x=62.5 That's your answer 62.5

opdaxy:

dang how yk this stuff

opdaxy:

i got 6 more for you if thats ok with u???

XioGonz:

It's easy when you use the simple rule of 3

XioGonz:

Yeah bring them in. Just remember to tag me

opdaxy:

dis 7th grade work

opdaxy:

ok

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz You need to occupy the simple rule of 3 in this case. CONGRATS @XioGonz Good job Scale on the map--2cm=25m(miles) For each 2cm=real distance of 25m(miles) Here we see " Dave measured the distance to the next town at 5 cm." So 2cm-25m(miles) 5cm-xm(miles) 2x=25*5 x=62.5 That's your answer 62.5 \(\color{#0cbb34}{\text{End of Quote}}\)

XioGonz:

Thank you!